OpenStudy (jkbo):

Please help me solve A sweater is on sale for 40% off the original price of $68. The sales tax rate is 7%. What is the total cost, including tax? A.45.56 B.43.66 C.47.60 D.40.80

OpenStudy (jdoe0001):

\(\bf n\% \ of \ m = \cfrac{ n }{ 100 }\times m\) what's 40% of $68?

OpenStudy (anonymous):

40/100 * 68 = 27.2 discount 68-27,2 = 40.8 sales tax = 7/100 * 68 = 4.76 total - 40.8 + 4.76 = 45.56

OpenStudy (jkbo):

Thank you both

OpenStudy (anonymous):

yw:D